Fama’s Llamas has a WACC of 9.6 percent. The company’s cost of equity is 11.8 percent, and its pretax cost of debt is 7.6 percent. The tax rate is 35 percent. What is the company’s target debt–equity ratio? (Do not round intermediate calculations and round your answer to 4 decimal places, e.g., 32.1616.) Debt–equity ratio
Answer #1

0.096 = w(0.076)(1 - 0.35) + (1 - w)(0.1180)

0.096 = 0.0494w + 0.1180 - 0.1180w

w = 0.3207

Weight of Debt = 0.3207

Weight of Equity = 1 - 0.3207 = 0.6793

Debt/Equity = 0.3207/0.6793

Debt/Equity = 0.47